Can someone please help me solve this 3x3 equation"
x-2y+3z=-4
3x+y-z=0
2x+3y+5z=1

Answer by richwmiller(17219)

x-2y+3z=-4
3x+y-z=0
2x+3y+5z=1
9x+3y-3z=0
x-2y+3z=-4
10x+y=-4
15x+5y-5z=0
2x+3y+5z=1
17x+8y=1
10x+y=-4
80x+8y=-32
17x+8y=1
63x=-33
x=-33/63
x=-11/21
-110/21+y=-4
-110/21+y=84/21
y=26/21
3x+y-z=0
-33/21+26/21-z=0
-7/21-z=0
-1/3=z
x = -11/21, y = 26/21, z = -1/3